# PentestAgent: Incorporating LLM Agents to Automated Penetration Testing

### v1

@@ 4.1.1. Benchmark Dataset
Several platforms can serve as the dataset of the benchmark, such as HackTheBox (HackTheBox, 2024), OWASP Benchmark (OWASP, 2024a), VulnHub (VulnHub, 2024), and VulHub (Vulhub, 2024). However, HackTheBox lacks accessibility to the public, requiring a VIP subscription to access most of its test machines, which creates a burden for using the benchmark. OWASP Benchmark and VulnHub contain thousands of target testing environments, covering a wide range of real-world penetration testing scenarios. However, setting up these environments for testing requires significant human effort. Furthermore, they do not provide a difficulty level reference for their test cases, necessitating manual effort to determine the difficulty level for each test case.
Finally, we chose VulHub as our benchmark dataset. VulHub provides an open-source collection of over a hundred pre-built vulnerable Docker environments, which has been widely recognized and utilized in penetration testing practices. The container-based platform supports infrastructure as code (IaC), making it easy to set up the testing environments. Besides, Docker containers provide sufficient isolation for penetration testing. Moreover, most vulnerable environments in VulHub are constructed to reproduce a particular Common Vulnerabilities and Exposures (CVE) (MITRE, 2024). Each vulnerable environment is associated with a CVE number, which allows us to use metrics associated with CVE numbers to learn about the properties of each vulnerable environment. Specifically, we learn about the difficulty of vulnerability exploits through the Common Vulnerability Scoring System (CVSS)(Database, 2024) and learn about how realistic the vulnerable environment is via the Exploit Prediction Scoring System (EPSS)(of Incident Response and Security Teams, 2024b). We elaborate on how we construct the benchmark dataset in §A.1 in the appendix.
As a result, we compiled a benchmark comprising 67 penetration testing targets, spanning 32 CWE (Common Weakness Enumeration) categories as shown in Fig.12 in the appendix. Within our benchmark, there are 50 targets with easy exploitability difficulty, 11 with medium exploitability difficulty, and 6 with hard exploitability difficulty. This diverse and realistic collection of vulnerable environments ensures a comprehensive assessment.
@@ 4.1.2. Metric
We measure the effectiveness of PentestAgent by determining whether all three stages of penetration testing are completed successfully and automatically. We define successful completion as follows: given a target IP, PentestAgent can automatically perform a functional exploit on the vulnerable environments.
Some penetration tests may be partially successful and require human assistance. However, failure in a previous penetration testing stage will affect the subsequent stages. To better understand the effectiveness of each component in PentestAgent, we measure the completion level at the stage level. This involves assessing the penetration testing stages that can be automatically completed, assuming the preceding stages have been successful. The completion criteria for each stage are defined as follows. The information gathering stage is considered complete if the target application is successfully identified by PentestAgent. The vulnerability analysis stage is marked as complete when PentestAgent identifies functional exploits based on the target application. We manually verify whether the discovered exploits are effective. The exploitation stage is completed if PentestAgent can automatically and successfully execute the exploit. This stage-level evaluation provides a granular understanding of PentestAgent ’s autonomy and effectiveness in progressing through the penetration testing process with minimal human assistance.
@@ 4.1.3. Environment setup
The simulated vulnerable applications are hosted on a virtual machine with 2 CPU cores and 8 GB RAM, running Ubuntu 22.04 LTS. To avoid interference with the testing process, we have disabled all services that require listening on ports, such as SSH.
The attacker machine is also hosted on a virtual machine with 16 CPU cores and 16 GB RAM, running Kali Linux 2024.1. The attacker machine includes all the pre-installed tools available in Kali Linux, with no additional tools installed.
The victim machine and the attacker machine maintain network connectivity via NAT. The vulnerable containers on the victim machine are created with the network parameter set to the victim machine’s IP, allowing the attacker machine to directly access the vulnerable environments hosted in the victim machine’s containers. This setup ensures the attacker can simulate real-world network conditions when attempting to exploit the vulnerabilities.
4.1.4. LLM models
We utilize the OpenAI GPT-3.5 and GPT-4 models, representing state-of-the-art LLM technology. These models were accessed via the OpenAI APIs. Specifically, we used the gpt-3.5-turbo-0125 model, which has a context window size of 16,385 tokens and is trained with data up to September 2021. The pricing for this GPT-3.5 model is \$0.50 / 1M input tokens and \$1.50 / 1M output tokens. Additionally, we used the GPT-4o model, which features a context window size of 128,000 tokens and is trained with data up to October 2023. The pricing for this GPT-4 model is \$5.00 / 1M input tokens and \$15.0 / 1M output tokens.
Our automated penetration testing framework does not extensively rely on LLMs’ inherent capabilities and knowledge to plan and perform most of the tasks. Instead, we designed a robust working pipeline for the penetration testing process, using LLMs as tools for specific and well-defined tasks, such as text summarization and code analysis. Given this design, we do not experiment with many other LLM models with our framework. This approach ensures consistency and reliability, as the framework’s performance is less dependent on the specific strengths and learned knowledge of different LLM models. By including GPT-3.5 and GPT-4, we aim to demonstrate the variability and invariability of our framework when utilizing models with varying capabilities, ensuring it is effective across different LLM configurations. In §5.2, we discuss the compatibility of PentestAgent with various LLMs.
@@ 4.1.4. LLM models
4.2. Effectiveness of the Entire Framework
4.3. Completion level of Penetration Testing Stages
@@
We further delve into PentestAgent by examining its completion level on individual penetration testing stages. Fig. 8 illustrates the completion levels of penetration testing stages across various difficulty categories and models. The PentestAgent with GPT-4 model exhibited robust performance across all stages and difficulty levels, demonstrating its ability to handle challenging tasks. In easy tasks, GPT-4 model achieved full completion in the intelligence gathering and vulnerability analysis stages, maintaining a high exploitation stage completion rate of 81.8%. In medium difficulty tasks, GPT-4 continued its strong performance in all stages, despite some drop in the vulnerability analysis stage. However, in harder tasks, the GPT-4 model encountered challenges, with completion rates dropping to 50% in the intelligence gathering stage, indicating potential limitations in handling complex scenarios that require better reconnaissance tools and advanced reasoning capabilities.
@@
Conversely, the GPT-3.5 model demonstrated varying degrees of success across different difficulty levels. It achieved very good performance in easier tasks, achieving high completion rates in intelligence gathering (92%) and vulnerability analysis (96%) stages, although with a slightly lower completion rate in exploitation (72%). In medium difficulty tasks, while maintaining a perfect record in vulnerability analysis (100%), it encountered difficulties in both intelligence gathering (66.7%) and exploitation (50%) stages, indicating challenges in complex reasoning required for effective reconnaissance and exploitation. Notably, in hard tasks, the GPT-3.5 model struggled significantly, achieving no completions in the intelligence gathering and exploitation stages, underscoring limitations in handling advanced penetration testing tasks that demand extensive contextual understanding and reasoning.
Overall, the findings suggest that while both models can automate substantial portions of penetration testing tasks, the GPT-4 model consistently outperforms the GPT-3.5 model, especially in more challenging scenarios.
Figure 9. Average time spent on penetration testing tasks
4.4. Efficiency
To evaluate PentestAgent’s efficiency, we measure the time spent and cost needed to perform penetration tests. Fig. 9 presents an overview of the average time and cost of conducting penetration testing tasks using the GPT-4 and GPT-3.5 models. Across all stages (intelligence gathering, vulnerability analysis, and exploitation), the GPT-4 model required, on average, 346.7 seconds for intelligence gathering, 780.9 seconds for vulnerability analysis, and 52.3 seconds for exploitation. The cumulative time spent on all tasks totaled 1164.7 seconds, with an average cost of \$2.66 per task. In contrast, the GPT-3.5 model exhibited lower time requirements and cost: 212.9 seconds for intelligence gathering, 698.8 seconds for vulnerability analysis, and 58.6 seconds for exploitation, resulting in a total average time of 1009.8 seconds and a cost of \$1.09 per task.
The analysis underscores the trade-off between efficiency and performance. While the GPT-4 model demonstrates superior capabilities in automated penetration testing, it does so at a higher operational cost.
@@ 4.5. Failure Analysis
4.6. Comparison with PentestGPT
We conducted a comparison of the effectiveness and efficiency of PentestAgent against PentestGPT. Unlike PentestAgent, PentestGPT requires human involvement for feedback and decision-making throughout the penetration testing process. Thus, we compare their performance using case studies. To create a fair evaluation, we randomly selected five vulnerabilities, including two categorized as easy, two as medium, and one as hard. We enlisted an undergraduate student with limited penetration testing experience to act as the human component required by PentestGPT. The student followed PentestGPT’s guidance without applying external knowledge for decision-making or task completion.
Under this testing condition, PentestGPT was unable to fully exploit any of the five vulnerabilities. In contrast, PentestAgent successfully completed exploitation in three out of the five cases. To provide a further comparison, we examined the performance of PentestGPT in individual penetration testing stages. PentestGPT was able to recognize the target application in only one of the five cases, whereas PentestAgent correctly identified the target application in four of the five cases. In the information gathering stage, PentestGPT spent an average of 826.25 seconds and required 7.4 rounds of interaction between the tester and the system. By comparison, PentestAgent completed information gathering in under 400 seconds on average, with no need for human-tester interaction. Given the correct target application information, PentestGPT successfully guided the exploitation of only one vulnerability. PentestAgent, on the other hand, automatically exploited four vulnerabilities, including the hard case, once the target application was identified.
These results demonstrate that PentestAgent significantly outperforms PentestGPT in both effectiveness and efficiency, accomplishing penetration testing tasks autonomously without requiring human assistance.
@@ 5. Discussion
5.1. Benchmark Coverage
5.2. Compatibility with Other LLMs
In light of data privacy concerns, PentestAgent has been designed to be compatible with a range of large language models (LLMs). Beyond commercial models like OpenAI’s pre-trained versions, we have also experimented with open-source LLMs such as Mistral and Llama 3. Our experiments confirm that PentestAgent functions effectively with SOTA open-source models, although the system’s efficiency can vary depending on the response time of the specific model in use. As we discussed in §4.5, the failure cases of PentestAgent are rarely caused by the capabilities of the LLM. Therefore, we believe that the overall effectiveness of PentestAgent is unlikely to be significantly impacted by the choice of LLM.
@@ 6. Conclusion
This paper presents PentestAgent, a novel LLM-based framework for automated penetration testing designed to address the limitations of existing frameworks: limited pentesting knowledge and insufficient automation. By leveraging a multi-agent architecture and incorporating various LLM techniques like Retrieval Augmented Generation (RAG), PentestAgent enhances the penetration testing process through improved knowledge integration and automation.
Our comprehensive benchmark, based on VulHub’s vulnerable Docker environments, provided a comprehensive test bed of PentestAgent. The evaluation results demonstrate that PentestAgent achieves satisfying performance in task completion and overall efficiency.
Appendix A Appendix

### v3

@@ 4.1.1. Benchmark Dataset
Several platforms can serve as the dataset of the benchmark, such as HackTheBox (HackTheBox, 2024), OWASP Benchmark (OWASP, 2024a), VulnHub (VulnHub, 2024), and VulHub (Vulhub, 2024). OWASP Benchmark and VulnHub contain thousands of target testing environments, covering a wide range of real-world penetration testing scenarios. However, setting up these environments for testing requires significant human effort. Furthermore, they do not provide a difficulty level reference for their test cases, necessitating manual effort to determine the difficulty level for each test case.
Finally, we chose VulHub and HackTheBox as our benchmark dataset. VulHub provides an open-source collection of over a hundred pre-built vulnerable Docker environments, which has been widely recognized and utilized in penetration testing practices. The container-based platform supports infrastructure as code (IaC), making it easy to set up the testing environments. Besides, Docker containers provide sufficient isolation for penetration testing. Moreover, most vulnerable environments in VulHub are constructed to reproduce a particular Common Vulnerabilities and Exposures (CVE) (MITRE, 2024). Each vulnerable environment is associated with a CVE number, which allows us to use metrics associated with CVE numbers to learn about the properties of each vulnerable environment. Specifically, we learn about the difficulty of vulnerability exploits through the Common Vulnerability Scoring System (CVSS)(Database, 2024) and learn about how realistic the vulnerable environment is via the Exploit Prediction Scoring System (EPSS)(of Incident Response and Security Teams, 2024b). We elaborate on how we construct the benchmark dataset in §B.1 in the appendix.
As a result, we compiled a benchmark comprising 67 penetration testing targets, spanning 32 CWE (Common Weakness Enumeration) categories as shown in Fig.13 in the appendix. These vulnerabilities cover eight security risks in OWASP Top 10 vulnerability (OWASP, 2024b). Within our benchmark, there are 50 targets with easy exploitability difficulty, 11 with medium exploitability difficulty, and 6 with hard exploitability difficulty. In addition, we incorporated 11 Capture The Flag (CTF) challenges from HackTheBox to simulate more challenging and realistic scenarios. These challenges are used in Section 4.5 for a practicality study and in Section 4.6 for the comparative evaluation with PentestGPT. This diverse and realistic collection of vulnerable environments ensures a comprehensive assessment.
@@ 4.1.2. Metric
We measure the effectiveness of PentestAgent by determining whether all three stages of penetration testing are completed successfully and automatically. We define successful completion as follows: given a target IP, PentestAgent can automatically perform a functional exploit on the vulnerable environments. For the HackTheBox targets, our focus is on obtaining initial access to the target host. In this context, a successful exploit is one that grants access to the target system.
Some penetration tests may be partially successful and require human assistance. However, failure in a previous penetration testing stage will affect the subsequent stages. To better understand the effectiveness of each component in PentestAgent, we measure the completion level at the stage level. This involves assessing the penetration testing stages that can be completed, assuming the preceding stages have been successful. The completion criteria for each stage are defined as follows. The information gathering stage is considered complete if the target application is successfully identified by PentestAgent. The vulnerability analysis stage is marked as complete when PentestAgent identifies functional exploits based on the target application. We manually verify whether the discovered exploits are effective. The exploitation stage is completed if PentestAgent can automatically and successfully execute the exploit. This stage-level evaluation provides a granular understanding of PentestAgent ’s autonomy and effectiveness in progressing through the penetration testing process with minimal human assistance.
@@
Table 2. Summary of LLM models used in our evaluation. Input/output costs are based on pricing at the time of testing.
Model Context Window Knowledge Cutoff Input Cost Output Cost
GPT-3.5-turbo-0125 16,385 tokens Sep 2021 \$0.50/1M tokens \$1.50/1M tokens
GPT-4o 128,000 tokens Oct 2023 \$5.00/1M tokens \$15.00/1M tokens
o1-mini 128,000 tokens Oct 2023 \$1.10/1M tokens \$4.40/1M tokens
Llama 3.1-8B-Instruct 128,000 tokens Dec 2023 Open-source Open-source
@@ 4.1.3. Environment setup
The simulated vulnerable applications are hosted on a virtual machine with 2 CPU cores and 8 GB RAM, running Ubuntu 22.04 LTS. To avoid interference with the testing process, we have disabled all services that require listening on ports, such as SSH. The attacker machine is also hosted on a virtual machine with 16 CPU cores and 16 GB RAM, running Kali Linux 2024.1. The attacker machine includes all the pre-installed tools available in Kali Linux, with no additional tools installed. The victim machine and the attacker machine maintain network connectivity via NAT. The vulnerable containers on the victim machine are created with the network parameter set to the victim machine’s IP, allowing the attacker machine to directly access the vulnerable environments hosted in the victim machine’s containers. This setup ensures the attacker can simulate real-world network conditions when attempting to exploit the vulnerabilities.
We evaluate our framework using a mix of commercial and open-source LLMs. Table 2 summarizes their key properties.
4.2. Effectiveness of the Entire Framework
@@ 4.3. Completion level of Penetration Testing Stages
To further evaluate PentestAgent, we analyze its performance across individual penetration testing stages. Fig. 8 presents the completion rates for intelligence gathering, vulnerability analysis, and exploitation across different difficulty levels and LLM backbones.
@@
GPT-4 demonstrated robust performance across all stages and difficulty levels, effectively handling a range of penetration testing tasks. In easy tasks, it achieved full completion in both intelligence gathering and vulnerability analysis, with an 81.8% completion rate in exploitation. For medium-difficulty tasks, GPT-4 maintained high completion rates across all stages, with a minor drop in vulnerability analysis. However, in hard tasks, performance declined, particularly in intelligence gathering (50%), suggesting limitations in handling complex reconnaissance tasks that require advanced reasoning and adaptive strategies.
@@
In contrast, GPT-3.5 exhibited more variability across difficulty levels. It performed well in easy tasks, with 92% completion in intelligence gathering and 96% in vulnerability analysis, though its exploitation stage completion rate (72%) was slightly lower. For medium tasks, while maintaining a 100% completion rate in vulnerability analysis, its performance declined in intelligence gathering (66.7%) and exploitation (50%), indicating challenges in navigating complex reconnaissance and execution scenarios. Notably, in hard tasks, GPT-3.5 struggled significantly, failing to complete both intelligence gathering and exploitation, highlighting its limitations in reasoning and contextual understanding required for advanced penetration testing.
Overall, both models effectively automate significant portions of penetration testing, but GPT-4 consistently outperforms GPT-3.5, particularly in more complex scenarios. These results emphasize the importance of advanced reasoning capabilities and enhanced reconnaissance strategies in achieving higher success rates in automated penetration testing.
(a) Completion level on targets
(b) Average time spent and cost on targets
Figure 9. Completion level and overhead of different LLM Backbones.
4.4. Ablation Study
To assess how different LLM backbones influence PentestAgent’s performance, we performed an ablation study on a subset of VulHub targets. Due to hardware limitations, specifically insufficient GPU resources to efficiently run the Llama 3.1 model on the full dataset, we selected a smaller subset comprising six easy, five medium, and two hard targets. Fig. 9(a) and Fig. 9(b) summarize the completion levels and overhead results, respectively. The results show that while all models perform consistently in vulnerability analysis (with 100% completion), differences emerge in the other stages. For the intelligence gathering stage, GPT-4 and o1-mini achieve higher completion levels compared to GPT-3.5 and Llama 3.1-8B-Instruct, suggesting that certain models are more effective in integrating context and managing complex reasoning. In the exploitation stage, o1-mini outperforms the others, indicating its strength in executing detailed attack procedures, whereas GPT-3.5 and Llama 3.1-8B-Instruct fall behind.
Overhead measurements further highlight trade-offs between processing time and cost. Although GPT-3.5 is faster in intelligence gathering and is the most cost-effective, its lower exploitation performance suggests a potential compromise in handling complex tasks. In contrast, while o1-mini delivers the best exploitation completion, it incurs longer processing times during vulnerability analysis. The Llama 3.1-8B-Instruct model, despite having no cost, suffers from significant time overhead and lower exploitation performance, which may limit its practical use.
4.5. Practicality Study
To evaluate PentestAgent’s effectiveness in realistic settings, we deployed it to solve HackTheBox challenges. Unlike standardized benchmarks, these challenges simulate dynamic, real-world penetration testing tasks by presenting diverse vulnerabilities across various CWEs. In this study, we selected 11 HackTheBox machines, nine labeled as easy, one as medium, and one as hard, to provide a broad assessment of the framework’s practical utility.
Fig. 10 shows the completion level and overhead of PentestAgent on HackTheBox challenges. Table 3 in Appendix C further details PentestAgent’s performance on these challenges by reporting the number of completed testing stages. While PentestAgent successfully exploited six machines, others like Pilgrimage achieved only partial completion, with only the vulnerability analysis stage being successful.
Overall, these results indicate that PentestAgent can address a range of real-world scenarios. However, the variability in stage completion underscores the need for further improvements to achieve more consistent, full-stage automation. We discuss the failed cases in detail in Section 4.7.
4.6. Comparison with PentestGPT
We conducted a comparison of the effectiveness and efficiency of PentestAgent against PentestGPT. Unlike PentestAgent, PentestGPT requires human participation for feedback and decision-making throughout the penetration testing process. Thus, we compare their performance using case studies. We randomly selected ten vulnerabilities from VulHub (five easy, three medium, and two hard) and included 11 HackTheBox challenges (nine easy, one medium, and one hard). Two evaluators with different skill levels conducted the tests: an undergraduate student with limited penetration testing experience evaluated PentestGPT on the VulHub targets, while a PhD student with more experience assessed the HackTheBox challenges. Both systems were configured to use the GPT-3.5 model to ensure a fair comparison.
Fig. 10 shows the completion level and overhead comparison between PentestAgent and PentestGPT on HackTheBox targets. Our results indicate that PentestAgent achieves higher exploitation success and overall efficiency. For instance, PentestAgent completes intelligence gathering in 220 seconds compared to 1199 seconds for PentestGPT, and finishes exploitation in 172 seconds versus 364 seconds. Although PentestGPT is slightly faster in vulnerability analysis, its slower performance in other stages due to human involvement reduces its overall efficiency. Additional results on VulHub targets and evaluation details can be found in the Appendix C. These findings imply that PentestAgent can deliver more consistent and timely penetration testing, highlighting the benefits of minimizing human intervention in real-world security assessments.
(a) Comparison of completion levels on HackTheBox targets
(b) Comparison of average time spent on HackTheBox targets
Figure 10. Completion level and overhead comparison on HackTheBox targets.
@@ 5. Discussion
5.1. Comparison with Existing Frameworks
While our primary comparison in this work is with PentestGPT, several emerging frameworks address related challenges in automated penetration testing. Notably, AutoAttacker and Enigma offer complementary perspectives that highlight different strengths and limitations relative to PentestAgent.
AutoAttacker (Xu et al., 2024) focuses exclusively on the post-breach stage of an attack. It is designed to automate the ”hands-on-keyboard” exploitation phase once a system has been compromised. In contrast, PentestAgent addresses the entire penetration testing pipeline—from reconnaissance and vulnerability analysis to exploitation—providing a more comprehensive solution. This broader scope is critical for real-world scenarios, where early-stage tasks are just as vital as post-breach actions for assessing and improving security.
Enigma (Abramovich et al., 2024), on the other hand, extends the SWE-agent framework by integrating interactive tools that support solving Capture The Flag challenges. Its design primarily targets challenges in the crypto and reverse engineering domains and still relies on human-in-the-loop interactions. Although Enigma’s interactive interfaces are effective for guiding the agent through specific problem domains, its reliance on manual intervention limits full automation. In contrast, PentestAgent is engineered to operate autonomously across diverse attack stages, reducing the need for human feedback and thus enabling more consistent performance in automated penetration testing.
Overall, these comparisons illustrate that while AutoAttacker and Enigma contribute valuable insights and capabilities, PentestAgent distinguishes itself by offering an end-to-end automated solution.
5.2. Limitations on Performing Sophisticated Pentesting
Our system, PentestAgent, focuses on exploiting individual vulnerable applications and services to help identify and mitigate these vulnerabilities. However, more sophisticated attack planning may be required in more complex penetration testing scenarios, such as red team simulations. These scenarios often involve combining several vulnerabilities to achieve a more challenging yet impactful exploit. For example, an SSRF vulnerability could be used as an intermediary step to exploit an internal application, eventually leading to obtaining root privileges.
While addressing such sophisticated attack strategies is beyond the scope of this paper, our framework, PentestAgent, can still be valuable in these complex scenarios. Our system can identify and validate exposed vulnerabilities, such as SSRF, which can serve as starting points for further exploitation. This initial identification and validation process can significantly contribute to the overall penetration testing workflow, providing a foundation upon which more advanced exploitation techniques can be built.
@@ 6. Conclusion
This paper presents PentestAgent, a novel LLM-based framework for automated penetration testing designed to address the limitations of existing frameworks: limited pentesting knowledge and insufficient automation. By leveraging a multi-agent architecture and incorporating various LLM techniques like retrieval augmented generation (RAG) and chain-of-thought (CoT), PentestAgent enhances the penetration testing process through improved knowledge integration and automation.
Our comprehensive benchmark, based on VulHub’s vulnerable Docker environments and HackTheBox CTF challenges, provided a comprehensive test bed of PentestAgent. The evaluation results demonstrate that PentestAgent achieves strong performance in task completion and overall efficiency.
Appendix A Prompts
This section specifies more prompts used in the PentestAgent pipeline.
The following prompt generates a structured output of the reconnaissance summary. Specifying the output structure and providing a comprehensive example guides the agent to output relevant information and reduces hallucination.
Reconnaissance Summary Prompt (Simplified)
  Provide a summary of all reconnaissance findings …
The summary of findings should be presented in valid JSON format with the following fields: {FORMAT SPEC.}
For example, {OUTPUT FORMAT EXAMPLE}
The following prompt summarizes the search results into a structured output for subsequent parsing and storing.
Search Results Summary Prompt
  List ALL CVE numbers, URLs, keywords, and their applicable version relevant to exploit the vulnerabilities of {APP}. The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
For example, {OUTPUT FORMAT EXAMPLE}
Exploit Procedure Analysis Prompt (Simplified)
RAG & CoT
Give a concise summary of the entire repository to answer the following questions:
1) whether this repository contains an exploit targeting a particular service or app;
2) What effect does the exploit have? Use one phrase to summarize the effect (e.g., remote command execution);
3) What relevant service/app version can this exploit be applied to?
Few-shot Learning
Note the app version is typically formatted as x.y.z. Explicitly state the version with the following formats …
4) what are the requirements to run this exploit? (e.g., OS, library dependencies, etc.)
Structured Output
You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …
For example, the response looks like this: {OUTPUT FORMAT EXAMPLE}
We designed the following prompt to generate a list of potential attack surfaces given a particular service or application.
Attack Surface Suggestion Prompt (Simplified)
List out all vulnerabilities ranked by confidence that can be used to exploit {app} {version} and provide the details about the vulnerabilities and the reasons to support each selection …
The details should include …
Make the selections by checking whether {version} is within the applicable version of the exploit and the vulnerability types …
The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
For example, {OUTPUT FORMAT EXAMPLE}
We designed the following prompt to generate a list of exploits for each potential attack surface.
Exploit Suggestion Prompt (Simplified)
List out paths of all relevant repositories ranked by the confidence that contain exploits …applicable to {app} {version} and provide the details about the exploit and reasons to support each selection …
The details should include …
Make the selections by checking whether {version} is within the applicable version of the exploit and the execution effects …
The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
For example, {OUTPUT FORMAT EXAMPLE}
Execution Information Query Prompt (Simplified)
  Based on the known information, try to provide the information listed here. {INFO NEEDED …}
CoT
You should examine the information needed one by one. For each piece of information needed, you should …
RAG
You should use your query tool to learn about the target environment …
Structured Output
The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
For example, the response looks like this: {OUTPUT FORMAT EXAMPLE}
Appendix B Benchmark Construction
B.1. VulHub Benchmark Construction
In addition, we remove the Docker images that are not associated with a CVE number and do not have CVSS 3.x scores. Additionally, some vulnerable applications are removed from the dataset due to complicated setup processes, such as requiring a license key from a service provider. To maintain integrity and fairness in our evaluations, we strictly prohibit PentestAgent from directly accessing any content from VulHub repository, thereby preventing any advantage or bias in our testing methodology.
Fig. 14 shows the difficulty rating distribution of our VulHub benchmark dataset.